Given the on-going pandemic, the sale of Willis cleaning

products jumped 15% to Rs. 425 million and increased income by 28%.

Household sales surged 7% and profits rose 65% as both prices and

demand increased and a germ-averse public stocked up. Assume yourself

as an ambitious manager of Willis cleaning products, who wants to use

these changes in the 4th quarter to calculate the own-price elasticity of

demand. You have the data on changes in price and changes in quantity

sold during the quarter.


1. What would be the main challenge they would face in calculating a

reliable estimate of elasticity?

2. Suppose for the 3rd quarter this challenge did not exist, rather the

challenge was one of data availability. All the manager knows, a few

minutes before a quickly called meeting to discuss 3rd quarter

results, is that the price of a pack of 35-count Clorox Wipes increased

by 12% while revenue from those wipes increased by 5%. Calculate

the own-price elasticity of demand for Clorox Wipes in this quarter.


1
Expert's answer
2022-02-23T12:22:12-0500

What would be the main challenge they would face in calculating a Reliable estimate of

elasticity?

The inability to capture the entire elasticity of their product when data is only available for one quarter is Willis Company's major hurdle in generating a realistic estimate of elasticity. According to the data, there has been an increase in the number of new germ-averse strains that have not been covered by the whole year, as well as an emergence of new germ-averse strains that have not been covered by the entire year. Given the present scenario and limited data, Willis Company may not be able to provide an accurate result in the fourth quarter. The unexpected surge in demand is caused by present conditions and it does not occur throughout the year.


2. Suppose for the 3rd quarter this challenge did not exist, rather the challenge was one of data availability. All the manager knows, a few minutes before a quickly called meeting to discuss 3rd quarter results, is that the price of a pack of 35-count Clorox Wipes increased by 12% while revenue from those wipes increased by 5%. Calculate the own-price elasticity of demand for Clorox Wipes in this quarter.

Own price elasticity is defined as "E_{QxPx}=\frac{\% ΔQ_x}{\%ΔP_x}"


Taking P as the price at the beginning of the quarter and P’ as the price at the end:

Rev = P × Q and P' = P×1.12 and R' = R×1.05

This means that:

"Q'=\frac{R'}{P'}" rearranging, we have;


"\frac{R}{P} \times \frac{1.05}{1.12} = Q \times \frac{1.05}{1.12}"

= Q × 0.0375 = Q × (1 - 0.0625)

To say, Q deceased by 6.25%

We can now calculate the elsaticity as:

"E_{QxPx}=\frac{\% ΔQ_x}{\%ΔP_x} = \frac{-0.0625}{0.12} = -0.0521"


Therefore, the own-price elasticity is -0.0521